Bob is gambling in the casino. The probability that he will win nothing is 0.2. The probability that he will

     win $10 is 0.3. The probability that he will win $20 is 0.5. What is the mean amount Bob will win?
Answer #1

from above probability distribution of Bob net gain:

P(X=0)=0.2

P(X=10)=0.3

P(X=20) =0.5

mean amount Bob will win E(x) =xP(x) =0*0.2+10*0.3+20*0.5 =13